OpenStudy (anonymous):

y=x+2 x=y-2 yes its a function

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You are correct. y = x+2 is a function and so is its inverse y = x-2 (since the original passes the horizontal line test)

OpenStudy (anonymous):

y=-x^2-3 = y=plus&minus sqr -y-3 and no its not a functon